# Reconstruction of gravitational form factors using generative machine learning

### v1

@@ V Summary and Outlook
We have developed a generative framework based on denoising diffusion for the model-independent reconstruction of hadronic form factors from sparse and noisy data. The training prior, built from ten distinct functional classes rooted in different theoretical approaches to hadron structure, encodes a broad manifold of physically plausible shapes without privileging any single parametrization. Applied to the proton gravitational form factors $A(t)$, $J(t)$, and $D(t)$, the framework yields non-parametric reconstructions consistent with lattice QCD and with standard parametric fits across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$. A systematic data-ablation protocol demonstrates that the reconstruction remains robust even when only one or two conditioning points are retained, confirming the strong constraining power of the physically informed prior. The densely sampled output of the diffusion model enables a direct extraction of the chiral low-energy constants $c_{8}=-4.6\pm 0.8$ and $c_{9}=-0.61\pm 0.19$, in agreement with the independent dispersive determination of Ref. Cao:2025dkv , and yields the nucleon $D$-term $D(0)=-4.3\pm 0.8$ at the physical pion mass.
@@
The precision of the present extraction is limited primarily by the use of the ChPT expression for $D(t)$ at $\mathcal{O}(p^{2})$. A natural next step is the construction of the complete generally covariant chiral Lagrangian up to and including $\mathcal{O}(p^{4})$ with explicit $\Delta(1232)$ degrees of freedom Alharazin:2023uhr . Matching the diffusion-model reconstruction to these higher-order expressions is expected to sharpen the determination of $c_{8}$ and $c_{9}$, reduce the uncertainty on the physical-mass extrapolation, and, in addition, provide access to the unknown LECs ($h_{i}$) of the $\Delta$-resonance Lagrangian appearing in Ref. Alharazin:2022wjj . This would yield the first model-independent, ChPT-based estimates of the GFFs and $D$-term of the $\Delta$ baryon. More broadly, the framework is general and can be applied to other hadronic form factors, including electromagnetic and transition form factors of baryons and mesons.
@@ Acknowledgements.
The authors thank Jambul Gegelia and Davide Laudicina for useful discussions. We are also grateful to Xiong-Hui Cao for clarifying the chiral order at which the low-energy constants in Ref. Cao:2025dkv were determined. The calculations of this work were performed on the HPC cluster Elysium of the Ruhr University Bochum, subsidised by the DFG (INST 213/1055-1).
@@
To verify robustness against the choice of conditioning strategy, we fine-tuned the converged model for 50 additional epochs at a flat learning rate of $2\times 10^{-5}$ ($5\times$ below the initial peak, to preserve the learned representations) and a broader five-type mask distribution. In addition to the three original mask types (see Sec. III.3), the expanded distribution introduces a high-$|t|$ cluster (5–15 points in the last 40% of the grid) and a sparse-spread mode (8–12 quasi-uniformly spaced points across the full grid). As shown in Fig. (3(b)), the validation loss, evaluated with the original mask distribution for comparability, remains within the 0.0170.0170.0210.021 band throughout, consistent with the initial convergence plateau. This confirms that the learned GFF manifold is stable and that reconstruction quality is insensitive to the specific mask distribution.
@@
(b) Fine-tuning with an expanded five-type mask distribution (epochs 101–150). The shaded band and dashed line show the validation range and best validation loss from the original training.
Abbildung 3: Training diagnostics for the $v$-prediction DDPM.
@@
where $\sigma_{i}$ represents the experimental uncertainty at each kinematic point. This procedure mimics realistic experimental conditions in which the true form factor is unknown and only noisy measurements at discrete momentum transfers are available. The diffusion model is then conditioned on the tuples $\{(t_{i},\,F^{\mathrm{data}}_{i},\,\sigma_{i})\}$ and tasked with reconstructing the full $t$-dependence, including extrapolation into unconstrained regions.
Starting from the full set of $N=5$ conditioning points, we progressively remove measurements from the high-$|t|$ end (rows 1–4 of Fig. (4)) and compare the resulting 68 % credible intervals against the known truth band. The bottom row provides the key comparison: three points retained at high $|t|$ instead of low $|t|$, with the same total number of constraints as row 3. Across all tests, the reconstruction degrades significantly when low-$|t|$ data, which anchor the form-factor normalisation, are removed, while high-$|t|$ points can be dropped with comparatively little loss of accuracy.
Abbildung 4: NN denotes the number of conditioning points. Rows 1–4 progressively remove high-$|t|$ conditioning points; row 5 keeps only high-$|t|$ points for contrast. The bottom row (red) uses the same number of points as the third row but retains only high-$|t|$ measurements, demonstrating the disproportionate importance of forward-region data.
Anhang D The proton GFFs at physical pion mass
Figure (5) presents the reconstruction of all three GFFs at the physical pion mass. For each form factor, the three panels show the effect of progressively reducing the number of ChPT conditioning points, from the full set (top) to a minimal subset (bottom). For $A(t)$ and $J(t)$ (Figs. (5(a)) and (5(b))), the Poincaré constraints at $t=0$ anchor the normalisation, and the reconstruction remains stable even as conditioning points are removed.
The form factor $D(t)$ (Fig. (5(c))) presents a qualitatively different situation. Unlike $A(t)$ and $J(t)$, no sum-rule constraint fixes $D(0)$, so the reconstruction relies entirely on the ChPT conditioning data. Moreover, the available ChPT expression for $D(t)$ is of lower chiral order than those for $A(t)$ and $J(t)$, so that the higher-$|t|$ conditioning points probe a regime where the truncated chiral expansion is no longer reliable and the quoted uncertainties underestimate the true error. The data-ablation sequence reveals that removing these higher-$|t|$ points actually improves the reconstruction: the diffusion-model median is not pulled toward the potentially unreliable high-$|t|$ ChPT values but instead follows a trajectory shaped by the training prior, effectively filtering out the incomplete chiral information.
$D(0)\;=\;c_{8}\,m_{N}\;=\;-4.3\;\pm\;0.8\,$, (18)
Abbildung 5: Diffusion-model reconstruction of the proton GFFs at the physical pion mass ($m_{\pi}\approx 139$ MeV), conditioned on ChPT data points evaluated with the LECs extracted in Sec. IV.3. Panel layout follows the data-ablation protocol: top panels use the most conditioning points, successive panels progressively remove them.
@@
  • (2) J. Y. Panteleeva, E. Epelbaum, J. Gegelia and U. G. Meißner, Eur. Phys. J. C 83 (2023) no.7, 617 [arXiv:2211.09596 [hep-ph]].
@@
  • (4) X. Ji and C. Yang, [arXiv:2508.16727 [hep-ph]].
@@
  • (25) L. Wang, G. Aarts and K. Zhou, JHEP 05 (2024), 060 doi:10.1007/JHEP05(2024)060 [arXiv:2309.17082 [hep-lat]].

### v2

@@ VI Summary and Outlook
We have developed a generative framework based on denoising diffusion for the model-independent reconstruction of hadronic form factors from sparse and noisy data. The training prior, built from ten distinct functional classes rooted in different theoretical approaches to hadron structure, encodes a broad manifold of physically plausible shapes without privileging any single parametrization. Applied to the proton gravitational form factors $A(t)$, $J(t)$, and $D(t)$, the framework yields non-parametric reconstructions consistent with lattice QCD and with standard parametric fits across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$. A systematic data-ablation protocol demonstrates that the reconstruction remains robust even when only one or two conditioning points are retained, confirming the strong constraining power of the physically informed prior. The densely sampled output of the diffusion model enables a direct extraction of the chiral low-energy constants $c_{8}=-4.6\pm 0.8\ \rm GeV^{-1}$ and $c_{9}=-0.61\pm 0.19\ \rm GeV^{-1}$, in agreement with the independent dispersive determination of Ref. Cao:2025dkv , and yields the nucleon $D$-term $D(0)=-4.3\pm 0.8$ at the physical pion mass.
@@
The precision of the present extraction is limited primarily by the use of the ChPT expression for $D(t)$ at $\mathcal{O}(p^{2})$. A natural next step is the construction of the complete generally covariant chiral Lagrangian up to and including $\mathcal{O}(p^{4})$ with explicit $\Delta(1232)$ degrees of freedom Alharazin:2023uhr . Matching the diffusion-model reconstruction to these higher-order expressions is expected to sharpen the determination of $c_{8}$ and $c_{9}$, reduce the uncertainty on the physical-mass extrapolation, and, in addition, provide access to the unknown LECs ($h_{i}$) of the $\Delta$-resonance Lagrangian appearing in Ref. Alharazin:2022wjj . This would yield the first model-independent, ChPT-based estimates of the GFFs and $D$-term of the $\Delta$ baryon.
As already mentioned, this approach can be applied to reconstruct the GFFs of various hadrons, such as $\Delta$-resonances, $\rho$-mesons, or even nuclear systems, including electromagnetic and transition form factors. Furthermore, our framework is naturally extensible to the reconstruction of GPDs, where the generative prior can be built from synthetic ensembles constructed via Double Distributions, ensuring that fundamental constraints such as polynomiality and positivity are satisfied by construction. Experimental Compton Form Factors (CFFs), extracted from DVCS measurements, then serve as the conditioning data, while GFF sum rules provide additional moment constraints to further regularize the reconstruction. Work along these lines is currently underway.
@@ Acknowledgements.
The authors thank Jambul Gegelia and Davide Laudicina for useful discussions. We are also grateful to Xiong-Hui Cao for clarifying the chiral order at which the low-energy constants in Ref. Cao:2025dkv were determined, and thank Xiangdong Ji and Latifa Elouadrhiri for helpful comments. The calculations of this work were performed on the HPC cluster Elysium of the Ruhr University Bochum, subsidised by the DFG (INST 213/1055-1).
@@
To verify robustness against the choice of conditioning strategy, we fine-tuned the converged model for 50 additional epochs at a flat learning rate of $2\times 10^{-5}$ ($5\times$ below the initial peak, to preserve the learned representations) and a broader five-type mask distribution. In addition to the three original mask types (see Sec. III.3), the expanded distribution introduces a high-$|t|$ cluster (5–15 points in the last 40% of the grid) and a sparse-spread mode (8–12 quasi-uniformly spaced points across the full grid). As shown in Fig. (7(b)), the validation loss, evaluated with the original mask distribution for comparability, remains within the 0.0170.0170.0210.021 band throughout, consistent with the initial convergence plateau. This confirms that the learned GFF manifold is stable and that reconstruction quality is insensitive to the specific mask distribution. We have further verified that varying the network architecture and training strategy leaves the validation loss within the same fluctuation band, indicating that systematic errors from these design choices are subdominant to the statistical noise floor.
@@
(b) Fine-tuning with an expanded five-type mask distribution. The shaded band and dashed line show the validation range and best validation loss from the original training.
Abbildung 7: Training diagnostics for the $v$-prediction DDPM.
@@
where $\sigma_{i}$ represents the experimental uncertainty at each kinematic point. This procedure mimics realistic experimental conditions in which the true form factor is unknown and only noisy measurements at discrete momentum transfers are available. The diffusion model is then conditioned on the tuples $\{(t_{i},\,F^{\mathrm{data}}_{i},\,\sigma_{i})\}$ and tasked with reconstructing the full $t$-dependence, including extrapolation into unconstrained regions. Starting from the full set of $N=5$ conditioning points, we progressively remove measurements from the high-$|t|$ end (rows 1–4 of Fig. (8)) and compare the resulting 68 % credible intervals against the known truth band. The bottom row provides the key comparison: three points retained at high $|t|$ instead of low $|t|$, with the same total number of constraints as row 3. Across all tests, the reconstruction degrades significantly when low-$|t|$ data, which anchor the form-factor normalisation, are removed, while high-$|t|$ points can be dropped with comparatively little loss of accuracy.
Abbildung 8: NN denotes the number of conditioning points. Rows 1–4 progressively remove high-$|t|$ conditioning points; row 5 keeps only high-$|t|$ points for contrast. The bottom row (red) uses the same number of points as the third row but retains only high-$|t|$ measurements, demonstrating the disproportionate importance of forward-region data.
@@
  • (2) J. Y. Panteleeva, E. Epelbaum, J. Gegelia and U.-G. Meißner, Eur. Phys. J. C 83 (2023) no.7, 617 [arXiv:2211.09596 [hep-ph]].
@@
  • (4) X. Ji and C. Yang, Nucl. Phys. B 1024 (2026), 117342 [arXiv:2508.16727 [hep-ph]].
@@
  • (10) V. D. Burkert, L. Elouadrhiri and F. X. Girod, Nature 557 (2018) no.7705, 396-399
  • (11) V. D. Burkert, L. Elouadrhiri and F. X. Girod, [arXiv:2104.02031 [nucl-ex]].
@@
  • (27) L. Wang, G. Aarts and K. Zhou, JHEP 05 (2024), 060 [arXiv:2309.17082 [hep-lat]].
@@
  • (66) Y. Guo, F. Yuan and W. Zhao, Phys. Rev. Lett. 135 (2025) no.11, 111902 [arXiv:2501.10532 [hep-ph]].
  • (67) K. Kumerički, Nature 570 (2019) no.7759, E1-E2.
  • (68) H. Dutrieux, C. Lorcé, H. Moutarde, P. Sznajder, A. Trawiński and J. Wagner, Eur. Phys. J. C 81 (2021) no.4, 300 [arXiv:2101.03855 [hep-ph]].